### v2

# Why Does New Knowledge Create Messy Ripple Effects in LLMs?

Abstract
Extensive previous research has focused on post-training knowledge editing (KE) for language models (LMs) to ensure that knowledge remains accurate and up-to-date. One desired property and open question in KE is to let edited LMs correctly handle ripple effects, where LM is expected to answer its logically related knowledge accurately. In this paper, we answer the question of why most KE methods still create messy ripple effects. We conduct extensive analysis and identify a salient indicator, GradSim, that effectively reveals when and why updated knowledge ripples in LMs. GradSim is computed by the cosine similarity between gradients of the original fact and its related knowledge. We observe a strong positive correlation between ripple effect performance and GradSim across different LMs, KE methods, and evaluation metrics. Further investigations into three counter-intuitive failure cases (Negation, Over-Ripple, Multi-Lingual) of ripple effects demonstrate that these failures are often associated with very low GradSim. This finding validates that GradSim is an effective indicator of when knowledge ripples in LMs.
[…]
Why Does New Knowledge Create Messy Ripple Effects in LLMs?
Jiaxin Qin1,2, Zixuan Zhang [^1] footnotemark: 1, Chi Han1, Manling Li1, Pengfei Yu1, Heng Ji1 1University of Illinois Urbana-Champaign 2Renmin University of China {qjx0814, zixuan11, chihan3, pengfei4, hengji}@illinois.edu manling@stanford.edu
1 Introduction and Related Work
Large language models (LLMs) can serve as powerful knowledge bases (KBs) thanks to their impressive knowledge storage, retrieval, and reasoning capabilities (Petroni et al., 2019; AlKhamissi et al., 2022). However, real-world knowledge keeps updating and evolving constantly, which motivates extensive research efforts on post-training knowledge editing (KE) Meng et al. (2022, 2023); Yin et al. (2023); Zhong et al. (2023); Song et al. (2024); Liu et al. (2024) to make sure that the knowledge in LMs remains accurate and up-to-date.
Figure 1: An illustration of ripple effects in LLM knowledge editing. Our work empirically demonstrates the positive correlation between gradient similarity explains a large portion of the ripple effect. Furthermore, messy similarities between knowledge points create several counter-intuitive ripple effect failures.
Previous research has proposed numerous evaluation metrics to ensure the efficiency and consistency of these editing methods. Among them, one critical criterion is the ability of the KE method to handle ripple effects Cohen et al. (2023), where a single edit should automatically and accurately propagate to related facts. For example, suppose an edit changes Leonardo DiCaprio’s nationality to Syrian. The model should automatically update its related information, such as knowing that his primarily used language is now Arabic. Such a task is very challenging because it requires the model to correctly understand and infer complex relationships among knowledge elements and accurately locate their parametric storage in order to perform the edits. Empirically, even though direct knowledge edits typically achieve over $>90\%$ accuracy, the success rates of ripple effects struggle to exceed $50\%$ across all recent KE methods, even on the simplest task in RippleEdits Cohen et al. (2023).
In this paper, we answer the intriguing research question of when and why updated knowledge ripples in language models. We hypothesize that the knowledge storage among parameters plays a critical role in determining the ripple effects between knowledge facts. A messy relationship among knowledge elements can make achieving a successful ripple effect intractable or impossible. Intuitively, the similarity of knowledge storage should be an important factor, as knowledge represented by similar parameters will respond similarly to parameter updates during knowledge editing. Following this intuition, we conduct extensive analysis and identify a salient indicator that strongly reveals how likely an updated fact will ripple in a language model: the cosine similarity between the gradients of the related knowledge facts (GradSim). We use gradients to represent knowledge storage distribution in LMs because they indicate which parameters in the model are responsible for increasing or decreasing the likelihood of answering certain knowledge. We observe a strong positive correlation between ripple effect performance and the cosine similarity of gradients across different LMs, editing methods, and evaluation metrics, with a Pearson correlation metric reaching as high as 0.85.
The hypothesis and analysis above predict a counter-intuitive phenomenon: knowledge with similar parameter-storing locations, even if logically unrelated or contradictory, will create positive ripple effects toward each other, and vise versa. Viewing GradSim as an indicator of ripple effects, we verify this paradox above by discovering and explaining three specific ripple effect cases: Negation, Over-Ripple Errors and Cross-Lingual Transfer. As illustrated in Figure 1, assuming that a knowledge edit changes the citizenship of Leonardo DiCaprio from United States to Syria. In Negation, LMs with different sizes including GPT2-XL and LLaMA-2 unexpectedly answer the negated query Leonardo is not a citizen of still by Syria instead of logically correct answers such as United States. Moreover, in Over-Ripple Errors, the LMs over-memorize the edit target Syria, and tends to always answer Syria even when asked about other topics such as language. In Cross-Lingual Transfer, even the most powerful cross-lingual LMs could make mistakes when asked about the edited knowledge in a different language. All of these ripple-effect cases are commonly encountered in real-world applications, but are more challenging and often experience counter-intuitive failures with current LMs and KE methods. In our experiments, we demonstrate that the model’s failure in these cases is strongly correlated with a too small GradSim, the similarity in knowledge distributions within LMs.
2 GradSim: A Ripple Effect Indicator
In this section, we formally introduce GradSim, a ripple-effect indicator based on the knowledge storage similarity between related knowledge. We use $x$ and $y$ to denote a pair of original fact and its related knowledge respectively, and we use $(q_{x},a_{x})$ and $(q_{y},a_{y})$ to represent query-answer pairs based on the corresponding knowledge facts. For instance, if $q_{x}$ and $a_{x}$ are <Leonardo DiCaprio is a citizen of> and <United States> respectively, then one example pair of $q_{y}$ and $a_{y}$ could be <Leonardo DiCaprio speaks> and <English>. Given a query $q_{x}$, typical KE methods update $a_{x}$ to a new answer $a^{\prime}_{x}$ by applying an update on the model parameters $\theta$, and ripple effect evaluations expect that the LM can automatically find the correct $a^{\prime}_{y}$ when asked $q_{y}$. Based on our hypothesis, knowledge represented by similar parameteres will respond similarly to parameter updates during knowledge editing. We employ the gradient to model the storage of knowledge within an LLM, and use the cosine similarity to measure the proximity between the storage distribution of two pieces of knowledge:
$\textit{GradSim}(x,y)=\cos(\nabla_{\theta}P_{\theta}(a_{x}|q_{x}),\nabla_{\theta}P_{\theta}(a_{y}|q_{y})).$
3 Experiments
We assess the effectiveness of GradSim by empirically examining its correlation with ripple effect performance, aiming to determine if it reliably indicates successful knowledge propagation in language models. Furthermore, we analyze three typical counter-intuitive failure cases in detail to understand the role of GradSim in these situations.
Figure 2: Main results of evaluating the correlation between ripple effect performances and GradSim values.
3.1 Data, Models, and KE Methods
In our experiments, we mainly employ RippleEdits Cohen et al. (2023), the most widely-used benchmark for evaluating ripple effects for knowledge editing methods in LLMs. We mainly use the popular split in RippleEdits because popular entities are more likely to be recognized by language models. This approach helps to minimize any side effects resulting from the model’s lack of knowledge, allowing us to focus on its reasoning abilities. To ensure a comprehensive evaluation, we also demonstrate results in the recent split as shown in Figure 2. In our experiments, we consider two typical knowledge editing (KE) methods: ROME Meng et al. (2022) and MEMIT Meng et al. (2023), which are based on the locate-and-edit approach to modify model parameters. For language models, we evaluate both larger models like LLaMA2-7B Touvron et al. (2023) and smaller models like GPT-2 XL Radford et al. (2019) to ensure a comprehensive evaluation and maintain consistency with the original settings in ROME and MEMIT.
3.2 Evaluation Metrics of Ripple Effects
To ensure the validity of our experiment results, we consider multiple evaluation metrics assessing how well the model performs in answering ripple-effect queries. These metrics include both accuracy-based measures, such as the exact match rate, and more quantified likelihood metrics, such as the absolute and relative gains in likelihood.
Exact-Match (EM) Rate
Similar to Cohen et al. (2023), we first consider accuracy-based metrics to calculate the proportion of correct answers the model generates from multiple random sampling choices. For each ripple query $q_{y}$, we sample 50 generated answers with a temperature 0.7, and compute the proportion of answers that include the correct answer. Our metric differs slightly from that in  Cohen et al. (2023), as we need to compute performance for each individual data point to analyze the overall correlation. The maximum length for generation is set to a small size of 15, as we believe that the answer is expected to appear early in a cloze-test query format.
Absolute Likelihood Gain
We also examine the answer probabilities to obtain a more detailed and quantifiable assessment of the performance. As probability values could be very small as the sequence length increases, we use log-likelihood score $\log P(a^{\prime}_{y}|q_{y})$, and measure its absolute gain on the correct answers before and after editing:
$\Delta\log P(y)=\log P_{\theta^{\prime}}(a^{\prime}_{y}\mid q_{y})-\log P_{\theta}(a^{\prime}_{y}\mid q_{y}).$
Relative Likelihood Gain
This metric is formulated by dividing the absolute gain of ripple effects by the absolute gain of the original fact, thereby normalizing the difficulty of the knowledge editing itself.
$\frac{\Delta\log P(y)}{\Delta\log P(x)}=\frac{\log P_{\theta^{\prime}}(a^{\prime}_{y}\mid q_{y})-\log P_{\theta}(a^{\prime}_{y}\mid q_{y})}{\log P_{\theta^{\prime}}(a^{\prime}_{x}\mid q_{x})-\log P_{\theta}(a^{\prime}_{x}\mid q_{x})}.$
3.3 Main Results
We conduct a comprehensive correlation analysis across different language models, knowledge editing methods, and performance metrics for ripple effects, with the results illustrated in Figure 2. A strong positive correlation is observed between ripple-effect performances and the GradSim values, validating that gradient-based knowledge storage similarity is a reliable indicator of ripple effects. Additionally, we observe the emergence of two distinct clusters in the figure. This clustering likely occurs because the data points can be categorized into successful and unsuccessful edits. Successful edits result in a significant improvement on performance, placing them in the upper successful cluster, while unsuccessful edits tends to remain in the unsuccessful cluster.
3.4 Counter-Intuitive Failure Cases
Negation
Negation is one of the most straight-forward ripple effects where the model is expected to answer a negated query after an editing is applied. For example, after editing the nationality of Leonardo DiCaprio as Syrian, the model should be able to avoid Syria given a negated query like “Leonardo DiCaprio is not a citizen of”. However, both smaller-sized LMs like GPT-2 and larger-sized LMs like LLaMA still answers “Syria” to this query and simply ignore the negation inside the sentence. In Figure 3, we first visualize both the values and gains of model likelihoods for the original and negated facts. The results demonstrate a strong positive (almost linear) correlation, indicating a severe problem of negation failures. In terms of GradSim values, we find that the gradient similarities between the original and negated facts are very high, suggesting that the original and negated facts are entangled in similar knowledge storage locations.
Figure 3: Correlation between original and negated facts on likelihood change, and the distributions of GradSim and likelihood ratios between original and negated facts.
Over-Ripple Errors
The over-ripple problem refers to the situation where, after a knowledge edit, the LM only memorizes the edited target itself and continues to provide this target as the answer even when asked about other knowledge that is related. For example, after editing the nationality of Leonardo DiCaprio as Syrian, the model will still answer Syria even when asked about the primary language Leonardo is speaking (the correct answer should be Arabic). In Figure 4, we first visualize GradSim distributions on $(q_{y},a^{\prime}_{x})$ and $(q_{y},a^{\prime}_{y})$ respectively, and we can observe that the edited target $a^{\prime}_{x}$ (e.g., Syria) has a much higher gradient similarity compared to the correct answer $a^{\prime}_{y}$ (e.g., Arabic). This explains the similar performance of answering both the correct and incorrect answers and indicates the occurrence of over-ripple errors.
Figure 4: The distributions of GradSim values and ripple effect performances.
Cross-Lingual Transfer
The problem of cross-lingual transfer is defined as the ability to edit a piece of knowledge in one language and have the model still provide the correct answer when asked a question in another language. We study the role of GradSim by visualizing the distribution of GradSim values and the ripple effect performance across different languages. We employ Baichuan Yang et al. (2023), the state-of-the-art bilingual model for Chinese and English. As shown in Figure 5, while the performance on the target language remains low, the GradSim values are also very low, primarily distributed near zero. GradSim works as a reliable indicator in this special ripple-effect case for cross-lingual transfer.
Figure 5: GradSim and performance distributions when editing on one language and testing on another.
4 Conclusion
Through extensive experiments and analysis, we propose GradSim, computed as the cosine similarity between gradients of the original fact and its related knowledge, as a crucial indicator for the effectiveness of the ripple effects. The positive correlation observed between GradSim and ripple effect performance across various LMs, KE methods, and evaluation metrics underscores its reliability. Additionally, our exploration of failure cases such as Negation, Over-Ripple, and Cross-Lingual scenarios, further confirms that low GradSim values are indicative of ripple effect failures.
Limitations
The first notable limitation is that, although a strong relationship between GradSim and the performance of ripple effects has been demonstrated, our research remains at the level of exploring correlations between these two factors and has not yet established a causal relationship. While it is always challenging to determine causality, it would still be extremely interesting and exciting to explore the dominant contributing factors to the complex distribution of knowledge storage in the pre-training phase of LMs. The second important limitation is that, while this paper identifies an indicator, we did not provide practical solutions for improving ripple effect performance by leveraging this indicator. However, we believe that the insights provided in this short paper will significantly enable the development of practical and effective methods in future research.
Appendix A Extra Results of GradSim
The RippleEdits dataset comprises six distinct tasks: Logical Generalization (LG), Compositionality I (CI), Compositionality II (CII), Subject Aliasing (SA), Preservation (PV), and Relation Specificity (RS) (Cohen et al., 2023). These tasks are designed to evaluate different aspects of knowledge editing in neural networks. Specifically, LG, CI, CII, and SA are tasks where the model is expected to manifest new knowledge in response to edits made to existing entries. Conversely, in the RS and PV tasks, the existing knowledge should remain unaltered post-editing, as these tasks are designed to test the model’s ability to preserve information that is logically independent from the changes applied.
In Figure 2, we analyze data from these tasks to illustrate a positive correlation between ripple effect performance and GradSim values across the four tasks explicitly associated with knowledge updates, hereafter referred to as ripple tasks. Each data point in Figure 6 is labeled to demonstrate the consistent applicability of the GradSim metric across the individual sub-tasks. In contrast, Figure 7 focuses on the two non-ripple tasks (RS and PV), where no significant correlation is observed between GradSim values and $\Delta\log P(y)$. The comparison underscores that GradSim is a critical metric for evaluating ripple effects, as it shows no significant impact in the non-ripple tasks, confirming its relevance specifically in contexts where knowledge modifications are expected.
Figure 7: Comparison of the Correlation on None-Ripple Tasks and Ripple Tasks.
Appendix B How to Represent Knowledge Distribution?
In this paper, we utilize gradients to represent the distribution of knowledge. To support this approach, we conducted preliminary experiments that lend credence to the underlying rationale of this intuition.
B.1 Does the way that we express a piece of knowledge change the knowledge distribution?
Figure 8: L1 Norm Distribution over LlaMA-7B
In this study, we calculate the gradient of a specific piece of knowledge, "The name of the currency in the country of citizenship of Leonardo DiCaprio is the Syrian pound," along with its variants: "The currency Leonardo DiCaprio uses is the Syrian pound" and "What’s the name of the currency in the country of citizenship of Leonardo DiCaprio? Syrian pound." We then plot the L1 norm of the gradient across the 32 downward projection layers of LlaMA7b. Prior research suggests that these layers are particularly adept at storing knowledge. Our results indicate that the distribution across these variants is remarkably consistent, suggesting that a piece of knowledge may be encoded within specific parameters of a large language model.